# Impact of Phonetics on Speaker Identity in Adversarial Voice Attack

### v1

@@ Abstract
Adversarial perturbations in speech pose a serious threat to automatic speech recognition (ASR) and speaker verification by introducing subtle waveform modifications that remain imperceptible to humans but can significantly alter system outputs. While targeted attacks on end-to-end ASR models have been widely studied, the phonetic basis of these perturbations and their effect on speaker identity remain underexplored. In this work, we analyze adversarial audio at the phonetic level and show that perturbations exploit systematic confusions such as vowel centralization and consonant substitutions. These distortions not only mislead transcription but also degrade phonetic cues critical for speaker verification, leading to identity drift. Using DeepSpeech as our ASR target, we generate targeted adversarial examples and evaluate their impact on speaker embeddings across genuine and impostor samples. Results across 16 phonetically diverse target phrases demonstrate that adversarial audio induces both transcription errors and identity drift, highlighting the need for phonetic-aware defenses to ensure the robustness of ASR and speaker recognition systems.
@@ I Introduction
Figure 1: Overview of the adversarial audio attack framework. A perturbation $\Delta x$ is applied to source audio to produce adversarial audio. The Automatic Speaker Recognition (ASR) model yields adversarial transcriptions, while the Speaker Identification (SID) model produces embeddings that drift from the source identity, degrading similarity.
While prior work has primarily characterized adversarial audio in terms of transcription metrics such as word error rate (WER) and character error rate (CER), comparatively little attention has been paid to the phonetic dimension of these perturbations. Phonetic units serve as the bridge between acoustic signals and linguistic representation; perturbations that alter phoneme realizations can not only change transcribed text but also modify salient cues used to characterize speaker identity. This dual effect suggests that adversarial perturbations may induce identity drift: a shift in the acoustic–phonetic footprint of a speaker that degrades the stability of downstream speaker recognition.
In this paper, we take a phonetic perspective on adversarial attacks against ASR. Using DeepSpeech [2] as our target model, we generate targeted adversarial examples and analyze their impact at two levels: phoneme-level confusions, and speaker identity embeddings. Our results show that adversarial perturbations often exploit systematic phonetic confusions, such as vowel centralization and consonant substitutions, which propagate into altered speaker embeddings. This not only produces misleading transcriptions but also erodes the consistency of speaker identity, demonstrating that adversarial audio attacks compromise ASR and speaker recognition simultaneously.
The contributions of this work are threefold: (i) we provide a phonetic-level analysis of adversarial audio, revealing how perturbations exploit systematic phoneme confusions; (ii) we utilize the concept of identity drift, characterizing how adversarial perturbations alter phonetic cues underlying speaker identity; and (iii) we conduct extensive experiments on DeepSpeech and evaluate speaker verification metrics (genuine vs. impostor samples), demonstrating the dual impact of adversarial perturbations.
@@ II Related Work
We evaluate the effect of adversarial perturbations on speaker identity using two state-of-the-art embedding models: ECAPA-TDNN [3] and a ResNet-based speaker verification backbone [4, 5]. These findings build on a large body of work showing that both ASR and speaker verification (SV) systems are vulnerable to adversarial manipulation. Hidden and practical command injection attacks such as CommanderSong [6], DolphinAttack [7], Hidden Voice Commands [8], CommanderGabble [9], and recent phoneme-level manipulations like PhantomSound [10] and SMACK [11] highlight real-world feasibility. Targeted adversarial examples against ASR were first demonstrated by Carlini and Wagner [1], with subsequent advances leveraging psychoacoustic models, masking, and over-the-air robustness [12, 13, 14]. More recent attack strategies include semantically meaningful perturbations [11], query-efficient black-box methods [10], and jailbreak-style attacks on audio-language models [15]. Speaker verification systems face similar threats: early works demonstrated targeted attacks against i-vector and DNN-based SV [16, 17, 18], with extensions to universal perturbations [9], robust real-time variants [17], and stress-testing embedding models [19, 20, 21]. Attacks exploiting voice conversion and spoofing further degrade reliability [22, 23], while studies such as Zhang et al. [24] demonstrate how silence and phonetic boundaries influence spoofing detection. On the defense side, Wu et al. [25] improved robustness via self-supervised pretraining, Chen et al. [26] leveraged codec bottlenecks for adversarial detection, and Kühne et al. [27] introduced diffusion models for ASR defenses. Complementary work includes DeepDetection for phishing prevention [28], vocal tract reconstruction for deepfake detection [23], and codec- or vocoder-based anti-spoofing pipelines [26, 24]. Despite this breadth, most approaches remain agnostic to the phonetic underpinnings of perturbations. Our work addresses this gap by explicitly connecting phoneme-level distortions—such as fricative instability, stop consonant confusion, and vowel centralization—to embedding drift, utilizing the concept of identity drift to unify transcription and speaker verification vulnerabilities.
@@ III-A Problem Formulation
Let $x\in\mathbb{R}^{T}$ denote an input audio waveform of length $T$. Our goal is to construct an adversarial example $x^{\prime}=x+\delta$ such that $x^{\prime}$ is perceptually indistinguishable from $x$ ($\|\delta\|$ is small under a distortion metric), yet simultaneously causes two effects: (i) the ASR system transcribes $x^{\prime}$ as a chosen target phrase $y_{t}$, and (ii) the speaker identity associated with $x$ drifts in the embedding space of a speaker identification (SID) system. This second phenomenon is known as identity drift. Formally, an attack succeeds if
@@ III-D Speaker Identity Evaluation
where $\mu_{\text{gen}}$ and $\mu_{\text{imp}}$ denote the means of the genuine and impostor score distributions, respectively, and $\sigma_{\text{gen}}^{2}$ and $\sigma_{\text{imp}}^{2}$ are their variances. A significant drop in $d^{\prime}$ indicates that adversarial perturbations shift embeddings closer to impostor space, thereby inducing identity drift. In addition, we report the true match rate at 0.1% false match rate (TMR@0.1%FMR). Using both ECAPA-TDNN and ResNet allows us to validate that observed identity drift is not model-specific but reflects vulnerabilities of embedding-based speaker recognition systems.
TABLE I: Representative 10 targets with short phonetic profiles and identity metrics per model. TMR measured at FMR $0.1\%$. $d^{\prime}$ from genuine vs. impostor score distributions.
@@ IV Experiments and Results
Figure 3: d’ across target texts (T1–T16) for ECAPA vs. RESNET50.
@@ VI Conclusion
Future research should, therefore, extend beyond controlled conditions to investigate phonetic-aware defenses, robustness training that explicitly accounts for phoneme-level vulnerabilities, and detection methods sensitive to the subtle score distribution shifts we identified. More broadly, adversarial attacks on audio warrant deeper study, particularly their interactions with linguistic variability, prosody, and channel effects. By bridging adversarial machine learning with phonetics, we take a step toward more secure and linguistically informed speech technologies. For reproducibility and extended visualizations, we provide additional figures in our GitHub repository. [^1] https://daniyalkabir.github.io/icassp-2025-results/

### v2

@@ Abstract
Adversarial perturbations in speech pose a serious threat to automatic speech recognition (ASR) and speaker verification by introducing subtle waveform modifications that remain imperceptible to humans but can significantly alter system outputs. While targeted attacks on end-to-end ASR models have been widely studied, the phonetic basis of these perturbations and their effect on speaker identity remain underexplored. In this work, we analyze adversarial audio at the phonetic level and show that perturbations are associated with systematic phonetic tendencies, such as vowel centralization and consonant substitutions. These distortions not only mislead transcription but also degrade phonetic cues critical for speaker verification, leading to identity drift. Using the DeepSpeech ASR model as our target, we generate targeted adversarial examples and evaluate their impact on speaker identity embeddings across genuine and impostor samples. Results across 16 phonetically diverse target phrases demonstrate that adversarial audio induces both transcription errors and identity drift, highlighting the need for phonetic-aware defenses to ensure the robustness of ASR and speaker recognition systems.
@@ I Introduction
Figure 1: Overview of the adversarial audio attack framework. A perturbation $\Delta x$ is applied to source audio to produce adversarial audio. The Automatic Speech Recognition (ASR) model yields adversarial transcriptions, while the Speaker Recognition model produces embeddings that drift from the source identity, degrading biometric similarity.
While prior work has primarily characterized adversarial audio in terms of transcription metrics such as word error rate (WER) and character error rate (CER), comparatively little attention has been paid to the phonetic dimension of these perturbations. Phonetic units serve as the bridge between acoustic signals and linguistic representation; perturbations that alter phoneme realizations can not only change transcribed text but also modify salient cues used to characterize speaker identity. This dual effect suggests that adversarial perturbations may induce identity drift: a shift in the acoustic–phonetic footprint of a speaker that degrades the stability of downstream biometric speaker recognition. While attack success is defined by forced transcription, identity drift is quantified post-attack via changes in speaker embedding similarity.
In this paper, we take a phonetic perspective on adversarial attacks against ASR. Using DeepSpeech [11] as our target model, we generate targeted adversarial examples and analyze their impact at two levels: phoneme-level confusions, and speaker identity embeddings. Our results show that adversarial perturbations are often associated with systematic phonetic confusions, such as vowel centralization and consonant substitutions, which propagate into altered speaker embeddings. This not only produces misleading transcriptions but also erodes the consistency of speaker identity, demonstrating that adversarial audio attacks compromise ASR and speaker recognition simultaneously.
The contributions of this work are threefold: (i) we provide a phonetic-level analysis of adversarial audio, revealing how perturbations exploit systematic phoneme confusions; (ii) we utilize the concept of identity drift, characterizing how adversarial perturbations alter phonetic cues underlying speaker identity; and (iii) we conduct extensive experiments on DeepSpeech and evaluate speaker recognition metrics, demonstrating the dual impact of adversarial perturbations.
@@ II Related Work
We evaluate the effect of adversarial perturbations on speaker identity using two state-of-the-art identity embedding models: ECAPA-TDNN [8] and a ResNet-based speaker verification backbone [12, 7]. These findings build on a large body of work showing that both ASR and speaker verification (SV) systems are vulnerable to adversarial manipulation. Hidden and practical command injection attacks such as CommanderSong [28], DolphinAttack [29], Hidden Voice Commands [3], CommanderGabble [31], and recent phoneme-level manipulations like PhantomSound [9] and SMACK [27] highlight real-world feasibility. Targeted adversarial examples against ASR were first demonstrated by Carlini and Wagner [4], with subsequent advances leveraging psychoacoustic models, masking, and over-the-air robustness [21, 20, 26]. More recent attack strategies include semantically meaningful perturbations [27], query-efficient black-box methods [9], and jailbreak-style attacks on audio-language models [13]. Localized and structured adversarial attacks have also been explored in the spectrogram domain, such as SpecPatch [10], which introduces a human-in-the-loop framework for crafting adversarial spectrogram patches against ASR systems. Speaker verification systems face similar threats: early works demonstrated targeted attacks against i-vector and DNN-based SV [18, 25, 5], with extensions to universal perturbations [31], robust real-time variants [25], and stress-testing embedding models [15, 17, 22]. Attacks exploiting voice conversion and spoofing further degrade reliability [2, 1], while studies such as Zhang et al. [30] demonstrate how silence and phonetic boundaries influence spoofing detection. On the defense side, Wu et al. [24] improved robustness via self-supervised pretraining, Chen et al. [6] leveraged codec bottlenecks for adversarial detection, and Kühne et al. [16] introduced diffusion models for ASR defenses. Complementary work includes DeepDetection for phishing prevention [14], vocal tract reconstruction for deepfake detection [1], and codec- or vocoder-based anti-spoofing pipelines [6, 30]. Despite this breadth, most approaches remain agnostic to the phonetic underpinnings of perturbations. Our work addresses this gap by explicitly connecting phoneme-level distortions—such as fricative instability, stop consonant confusion, and vowel centralization—to embedding drift, utilizing the concept of identity drift to unify transcription and speaker verification vulnerabilities.
@@ III-A Problem Formulation
Let $x\in\mathbb{R}^{T}$ denote an input audio waveform of length $T$. Our goal is to construct an adversarial example $x^{\prime}=x+\delta$ such that $x^{\prime}$ is perceptually indistinguishable from $x$ ($\|\delta\|$ is small under a distortion metric), yet simultaneously causes two effects: (i) the ASR system transcribes $x^{\prime}$ as a chosen target phrase $y_{t}$, and (ii) the speaker identity associated with $x$ drifts in the embedding space of a speaker recognition system. This second phenomenon is known as identity drift. Formally, an attack succeeds if
@@ III-D Speaker Identity Evaluation
where $\mu_{\text{gen}}$ and $\mu_{\text{imp}}$ denote the means of the genuine and impostor score distributions, respectively, and $\sigma_{\text{gen}}^{2}$ and $\sigma_{\text{imp}}^{2}$ are their variances. A significant drop in $d^{\prime}$ indicates that adversarial perturbations shift embeddings closer to impostor space, thereby inducing identity drift. In addition, we report the TMR at 0.1% FMR. Using both ECAPA-TDNN and ResNet allows us to validate that observed identity drift is not model-specific but reflects vulnerabilities of embedding-based speaker recognition systems.
TABLE I: Representative 10 targets with short phonetic profiles and identity metrics per model: TMR measured at FMR $0.1\%$ and $d^{\prime}$ from genuine vs. impostor score distributions.
@@ IV Experiments and Results
Figure 3: $d^{\prime}$ across target texts (T1–T16) for ECAPA vs. RESNET50.
@@ VI Conclusion
Future research should, therefore, extend beyond controlled conditions to investigate phonetic-aware defenses, robustness training that explicitly accounts for phoneme-level vulnerabilities, and detection methods sensitive to the subtle score distribution shifts we identified. More broadly, adversarial attacks on audio warrant deeper study, particularly their interactions with linguistic variability, prosody, and channel effects. By bridging adversarial machine learning with phonetics, we take a step toward more secure and linguistically informed speech technologies. For reproducibility and extended visualizations, we provide additional figures in our GitHub repository [^1] https://daniyalkabir.github.io/icassp-2026-results/.